OpenStudy (jenniferjuice):

Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (jenniferjuice):

4x + 3y = -1 +3x + y = 3 ---------------- 7x + 4y = 2 - 4y -4y --------------- 7x = 2 - 4y

OpenStudy (jenniferjuice):

ive got this now what do i do?

OpenStudy (anonymous):

You're supposed to make one of the variables disappear first: 4x + 3y = -1 (Line 1) +3x + y = 3 (Line 2) Multiply Line 2 by -3, giving -9x-3y=-9 Now add them together, and you get: 4x+(-9x)+3y+(-3y) = -1+(-9) -5x+0y = -10 x = 2. Now put x back into the original Line 2, and you will get: 3*(2)+y = 3 6 + y = 3 y = -3 And there you have it